OpenStudy (hba):

Value of x in log5x = 4logx5 is

OpenStudy (hba):

@hartnn

hartnn (hartnn):

draw the question

OpenStudy (hba):

Ok

OpenStudy (hba):

Ok\[\log_{5}x=4\log_{x}5\]

hartnn (hartnn):

use : \(\huge\log_ab= \frac{log b}{loga}\)

OpenStudy (hba):

logx/log5=4log5/logx whats next ??

hartnn (hartnn):

try to bring log x on one side and other terms on other side of = sign

OpenStudy (hba):

(logx)(logx)=4(log5)(log5)

OpenStudy (hba):

What do i do next ?? @hartnn

hartnn (hartnn):

yup, take square root on both sides after writing log x . logx as (log x)^2

OpenStudy (hba):

logx=2log5

hartnn (hartnn):

now use \(nlog x =log x^n\)

OpenStudy (hba):

logx=log5^2

hartnn (hartnn):

what is 5^2 ?

OpenStudy (hba):

25

OpenStudy (hba):

logx=log25

OpenStudy (hba):

so x=25

hartnn (hartnn):

yes..

OpenStudy (anonymous):

5 ^4

OpenStudy (hba):

but i think i will get 2 answers

OpenStudy (anonymous):

\[\log_{x}5^a = a \log_{x}5 \]

hartnn (hartnn):

how 2 ?

OpenStudy (anonymous):

use this so u will get x = 5^4 =625 , do u get it @hba

OpenStudy (hba):

25 and 1/25

OpenStudy (hba):

@akash_809 Please dont confuse me

OpenStudy (hba):

Options given .A) 25, 1/25 B) 1, -1 C) 0, 1 D) 2, 3 E) 4, 5

hartnn (hartnn):

when u take square root, u also get -*minus* so u also get 1/25

OpenStudy (hba):

Ok Boss Thanks :)

hartnn (hartnn):

welcomes :)